Reliability of the Models and Their Match with Observations

Steven E. Brown^{∗} and Gerard M. Steiner^{†}

∗National Center for Atmospheric Research^{1} Boulder Colorado 80307 †National Standards Institute, Boulder, Colorado 11543

Abstract. This is samp2.tex — a sample input file for aipproc. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system.

INTRODUCTION

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that stateof-the-model climate models are credible simulators of the true climate system [1–3]. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system.

ΔT_{2c} = \frac{S_{C}}{S_{T}} ln(2)  (1)

∑ |M_{g}^{viol}|^{2} = g_{S}^{2n−4}(Q^{2}) N^{n−2}(N^{2} − 1)

1) The National Center for Atmospheric Research is sponsored by the National Science Foundation.

× (∑_{i<j}) ∑_{perm} \frac{1}{S_{12}} \frac{1}{S_{12}} ∑_{τ} c_{τ}^{f} .  (2)

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system.

ΔT_{2c} = ∫_{0}^{t} F (τ) [\frac{S_{C}}{S_{T}} ] − S_{T} ΔT (t)  (3)

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that stateof-the-model climate models are credible simulators of the true climate system. See Figure 1.

SPECIAL NUMBERING AND EXTRA SYMBOLS

To obtain numbers not normally produced by the automatic numbering, use the \eqnum{#1} command, where #1 is the desired equation number. For example, to get an equation number of (1.2′),

FIGURE 1. Concentration profile climatology and climate model situation.

g^{+}g^{+} → g^{+}g^{+}g^{+}g^{+} …, q^{+}q^{+} → q^{+}g^{+}g^{+} ….  (1.2′)

A few notes on \eqnum{#1}. The \eqnum{#1} must come before the \label{#1}, if any. The numbering set with \eqnum{#1} is transparent to the automatic numbering in REVTEX; therefore, you must know the number ahead of time, and must make sure that the number set with \eqnum{#1} stays in step with the automatic numbering. \eqnum{#1} works with both single-line and multiline equations. You could, if you wished, do all the numbering in a paper manually with \eqnum{#1}.

Enclosing single-line and multiline equations in \begin{mathletters} and \end{mathletters} will produce a set of equations that are “numbered” with letters, as shown in Eqs. (4a) and (4b) below:

\{abc123456abcdefαβγδ1234556αβ \frac{1 ∑_{b}^{a}}{A^{2}} \} ,  (4a)

M = ig_{Z}^{2}(4E_{1}E_{2})^{1/2}(l_{i}^{2})^{−1}(g_{σ_{2}}^{e})^{2}χ_{−σ_{2}}(p_{2})
×[ϵ_{i}]_{σ_{1}}χ_{σ_{1}}(p_{1}).  (4b)

If you use a \label{#1} command right after the \begin{mathletters}, then \ref{#1} can be used to reference all the equations in a mathletters environment. For example, the equations in the preceding mathletters environment were Eqs. (4).

The AIP book style will support all the “normal” TEX and L^{A}TEX symbols, and in addition, the following symbols

\corresponds  =^{△}  \gtrsim     > ∼
\lesssim         < ∼  \precsim ≺_{∼}
\succsim  ≻_{∼}  \lambdabar λ

\tensor{T}       T^{↔}

The following is a longtable; See Table 1.

1

[Representative symbiotic stars.The organization of the archive was de bated at great length. We hope it will be useful. If you don’t like our organization, feel free to move the fil es around as you see fit, not forgetting to redefine the]Representative symbiotic stars.The organization of the archive was de bated at great length. We hope it will be useful. If you don’t like our organization, feel free to move the fil es around as you see fit, not forgetting to redefine the

Name  Other Name α_{2000} δ_{2000} l_{II} b_{II}
EG And  HD 4174 00:44:37.1 +40:40:46.7 121.5 -22.2
AX Per  MWC 411 01:36:17 +54:15:36 129.5 -8.0
BX Mon  AS 150 07:25:24  -03:36:00 220.0  5.9
RX Pup  HD 69190 08:14:12  -41:42:24 258.5 -3.9
SY Mus  HD 100336 11:32:05.5  -65:25:07 294.8 -3.8
AG Dra  – 16:01:40.9 +66:48:09.7 100.3 41.0
RT Ser  MWC 265 17:39:52  -11:56:42 13.9 10.0
AR Pav  MWC 600 18:20:28  -66:04:48 328.5 -21.6
BF Cyg  MWC 315 19:23:53.2 +29:40:26 62.9  6.7
CH Cyg  HD 182917 19:24:33.0 +50:14:29.7 81.8 15.6

[Continued.]Continued.

Name  Other Name α_{2000} δ_{2000} l_{II} b_{II}
HM Sge  – 19:41:57.0 +16:44:40 53.6 -3.2
CI Cyg  MWC 415 19:50:12 +35:41:06 70.9  4.7
V1016 Cyg  AS 373 19:57:05 +39:49:48 75.2  5.7
RR Tel  Hen 1811 20:04:16  -55:43:18 342.2 -32.2
V1329 Cyg  HBV 475 20:51:01 +35:34:54 77.8 -5.5
AG Peg  HD^{a} 207757 21:51:01.9 +12:37:32.0 69.3 -30.9
Z And^{b}  HD 221650 23:33:39.9 +48:49:06.3 110.0 -12.1
R Aqr  HD 222800 23:43:49.3 -15:17:02.7 66.5 -70.3

a Longtable note 1. b Longtable note 2.

POLICY-ANALYTIC APPROACH

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system. See Table 2.

Atmospheric Model

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system [2]. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that stateof-the-model climate models are credible simulators of the true climate system. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system.

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system.

TABLE 2. This table does not expand to the table does not ah ah ah expand to ta ble does NOT expand to textwidth.

Main sample  Secondary result  Previous result^{a}  Change^{b}(%)
Harkonen et al.  34.5  24.5  −23
Deckard and Batty  24.5  28.5  −13
Marcin  134.5  999999224.5  53

a First demonstrate that state-of-the-model climate models are credible simulators of the true climate system. b Estimated 2.

Quantitative Differences

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system.

High-T_{c} calculations. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system.

High-T_{c} calculations. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true dynamic climate system [3].

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true dynamic climate system.

Qualitative Differences

The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that state-of-the-model climate models are credible simulators of the true climate system. The objectives of my talk are to compare climate model situations with observed climatology and to demonstrate that stateof-the-model climate models are credible simulators of the true climate system.

REFERENCES

1. Smith A., and Doe B., J. Chem. Phys. 76, 4056 (1982).

2. Jones, C., J. Chem. Phys. 68, 5298 (1978).

3. Jones, C., and Smith, A., Title of Book, New York: IEEE Press, 1978, ch. 6, pp. 23-26.